# Fermi-Energy-Sensitive Universal Conductance Fluctuations in Anisotropic Materials

### v2

@@ III.2 Numerical results of UCF for the anisotropic free electron gas on a 2D lattice
In this subsection, we compare the numerical results of UCF from ensemble-averaged calculations to test the validity of our theoretical results. Throughout this paper, we fix the infinite left and right leads to have isotropic dispersion to avoid the influence of lead anisotropy, with the transport direction along $z$ unless otherwise stated. We consider the Hamiltonian $H^{cen}$ on a 2D lattice to lower the calculation cost.
@@
The conductance fluctuation $\Delta G$ as a function of disorder strength $W$ is shown in Fig. 3. With the increase of W, transport in the central region experiences transitions from a ballistic regime to a diffusive regime. Then the conductance vanishes after a metal-insulator transition for large enough $W$. In the diffusive regime, the conductance fluctuation plateau versus $W$ is a signature of the disorder-independent UCF and is used as a prescription to determine the amplitude of UCF. For example, in Fig. 3(a), we take the average value on the plateau of the $\Delta G$ vs $W$ trace to be the UCF amplitude. This value is close to the theoretical result of $0.390$ (marked by the horizontal dashed line) calculated using the formula in Appendix  A, where we have used $\Delta G=\tilde{c}_{d}\sqrt{\frac{ks^{2}}{\beta}}$ and the symmetry indices $k=1,\beta=1,s=1$ for the orthogonal ensemble under consideration. We have tested several sizes of $L_{y}=L_{z}=40,80,120,300$ to ensure that the influence of the finite-size effect is negligible. The UCF amplitude converges quickly to the theoretical value when the system size reaches $80\times 80$. Similar agreements have been observed for various degrees of band anisotropy, which is modulated by the anisotropy parameters $\frac{M_{z}}{M_{y}}=1.0,2.0,4.0$ in Fig. 3 (b-d). The overall increase of UCF with the increasing of $\frac{M_{z}}{M_{y}}$ is also consistent with our analysis in the subsection III.1.
Figure 3: (Color online) Numerical results of the conductance fluctuations $\Delta G$ versus the disorder strength W for different anisotropy parameters $\frac{M_{z}}{M_{y}}$ in the 2D (a-d) square system with $L_{y}/L_{z}=1,L_{x}=1$ and (e-h) rectangular system with $L_{y}/L_{z}=2,L_{x}=1$. (a) $\frac{M_{z}}{M_{y}}=0.4$. (b) $\frac{M_{z}}{M_{y}}=1.0$. (c) $\frac{M_{z}}{M_{y}}=2.0$. (d) $\frac{M_{z}}{M_{y}}=4.0$. The Horizontal dashed lines in (a-d) correspond to the theoretical values of 0.390, 0.431, 0.481, and 0.547, respectively. (e) $\frac{M_{z}}{M_{y}}=0.4$. (f) $\frac{M_{z}}{M_{y}}=1.0$. (g) $\frac{M_{z}}{M_{y}}=2.0$. (h) $\frac{M_{z}}{M_{y}}=4.0$. The Horizontal dashed lines in (e-h) correspond to the theoretical values of 0.463, 0.547, 0.628, and 0.728, respectively. The open boundary conditions are applied in the numerical calculations. The Fermi energy is $E_{F}=0.4$ in the central region and $E_{F}=0.6$ in the lead. The transport direction is along z, and the ensembles for average are 1200.
@@ V.1 Tight binding model of $\mathrm{Cd_{3}As_{2}}$
where $a^{\prime}_{\mathbf{r}}(a^{\prime\dagger}_{\mathbf{r}})$ represents the two-component electron annihilation(creation) operators on the site $\mathbf{r}$. Based on the parameters of the real material given in [78], we set the parameters as $C_{0}=-0.0145~{}eV,C_{1}=10.59~{}eV\overset{\circ}{A^{2}},C_{2}=11.5~{}eV\overset{\circ}{A^{2}}$, $M_{0}=0.0205~{}eV,M_{1}=-18.77~{}eV\overset{\circ}{A^{2}},M_{2}=-13.5~{}eV\overset{\circ}{A^{2}}$, the SOC strength $A=0.889~{}eV\overset{\circ}{A}$, and the lattice constant along the three directions is $a=b=3~{}\overset{\circ}{A},c=5~{}\overset{\circ}{A}$. These parameters apply to both the central region and the two leads. The Fermi energy in the lead is set equal to that in the central region to ensure state alignment. Open boundary conditions are applied in the conductance calculations. The onsite disorder is added and averaged over at least 1200 ensembles to calculate $\Delta G$. In the following, $eV$ and $\overset{\circ}{A}$ are taken as units of energy and length, respectively.
@@ V.2 UCF versus Fermi energy along y and z transport directions
The largest system size considered is $L_{x}=L_{y}=L_{z}=20$, where $L_{\alpha}$ counts the number of unit cells along the $\alpha$ direction. Unlike in previous sections, the lattice constants are not unity. The system is thus spatially anisotropic with lengths $\hat{L}_{x}=\hat{L}_{y}=60~{}\overset{\circ}{A}$ and $\hat{L}_{z}=100~{}\overset{\circ}{A}$, where $\hat{L}_{\alpha}$ measures the sample lengths in units of $\overset{\circ}{A}$ along the $\alpha$ direction. As discussed at the end of Section  IV.2, this geometry will favor a significant Fermi energy dependence of UCF for transport along the $x$ and $y$ directions by decreasing the diffusion time $t_{x/y}$ in Eq. 8 and will be demonstrated below.
@@ VI Discussions and Conclusions
Motivated by recent experimental observations [51] and theory progress [15], this work comprehensively studies UCF in materials with anisotropic band dispersion. This introduces a significant Fermi-energy dependence of UCF, in contrast to the Altshuler-Lee-Stone theory [1, 2, 7], whose quantitative result suggests a Fermi-energy independence.
@@
As an application, we calculate the conductance fluctuations in $\mathrm{Cd_{3}As_{2}}$ using first-principle parameters near the Dirac point. We consider a specific sample setup cleaved with the same number of unit cells along the $L_{\alpha}$($\alpha=x,y,z$) direction. It is found that the UCF increases when chemical potential moves away from the Dirac point for transport along the $x$ or $y$ direction. In contrast, for transport along the $z$ directions, the UCF dependence on Fermi energy is much less sensitive above the Dirac point. The distinct behaviors of UCFs are testable in experiments. The scope of this work addresses the Fermi energy dependence of UCF in anisotropic materials. Our work does not consider the effects of the magnetic field [85], the long-range impurities [38], decoherence [50, 86, 87, 57, 46], finite temperature [88, 89], the changes of symmetry indices [39, 40, 54], disorder-induced phase transitions [55, 30], non-ergodicity [56], etc, which will be left for future research. In addition to $\mathrm{Cd_{3}As_{2}}$, anisotropic materials are prevalent [72] in nature and can be engineered by straining [90, 91] or twisting  [92]. With the well-developed techniques of tuning the chemical potential through gate voltage or doping, we expect our findings to be readily testable in future experiments.

### v3

@@ III.2 Numerical results of UCF for the anisotropic free electron gas on a 2D lattice
In this subsection, we compare the numerical results of UCF from ensemble-averaged calculations to test the validity of our theoretical results. Throughout this paper, we fix the infinite left and right leads to have isotropic dispersion to avoid the influence of lead anisotropy, with the transport direction along $z$ unless otherwise stated. The UCF theory applies in any dimensions [22, 23]; however, to lower the calculation cost, we only perform detailed calculations for Hamiltonian $H^{cen}$ on a 2D lattice in this subsection. (A representative result showing comparisons between theory and numerics in 3D lattices has also been given in Appendix C.)
@@
The conductance fluctuation $\Delta G$ as a function of disorder strength $W$ is shown in Fig. 3. With the increase of W, transport in the central region experiences transitions from a ballistic regime to a diffusive regime. Then the conductance and its fluctuation eventually vanish after a metal-insulator transition for large enough $W$. In the diffusive regime, the conductance fluctuation plateau versus $W$ is a signature of the disorder-independent UCF and is used as a prescription to determine the amplitude of UCF. For example, in Fig. 3(a), we take the average value on the plateau of the $\Delta G$ vs $W$ trace to be the UCF amplitude. This value is close to the theoretical result of $0.390$ (marked by the horizontal dashed line) calculated using the formula in Appendix  A, where we have used $\Delta G=\tilde{c}_{d}\sqrt{\frac{ks^{2}}{\beta}}$ and the symmetry indices $k=1,\beta=1,s=1$ for the orthogonal ensemble under consideration. We have tested several sizes of $L_{y}=L_{z}=40,80,120,300$ to ensure that the influence of the finite-size effect is negligible. The UCF amplitude converges quickly to the theoretical value when the system size reaches $80\times 80$. Similar agreements have been observed for various degrees of band anisotropy, which is modulated by the anisotropy parameters $\frac{M_{z}}{M_{y}}=1.0,2.0,4.0$ in Fig. 3 (b-d). The overall increase of UCF with the increasing of $\frac{M_{z}}{M_{y}}$ is also consistent with our analysis in the subsection III.1.
Figure 3: (Color online) Numerical results of the conductance fluctuations $\Delta G$ versus the disorder strength W for different anisotropy parameters $\frac{M_{z}}{M_{y}}$ in the 2D (a-d) square system with $L_{y}/L_{z}=1,L_{x}=1$ and (e-h) rectangular system with $L_{y}/L_{z}=2,L_{x}=1$. (a) $\frac{M_{z}}{M_{y}}=0.4$. (b) $\frac{M_{z}}{M_{y}}=1.0$. (c) $\frac{M_{z}}{M_{y}}=2.0$. (d) $\frac{M_{z}}{M_{y}}=4.0$. The Horizontal dashed lines in (a-d) correspond to the theoretical values of 0.390, 0.431, 0.481, and 0.547, respectively. (e) $\frac{M_{z}}{M_{y}}=0.4$. (f) $\frac{M_{z}}{M_{y}}=1.0$. (g) $\frac{M_{z}}{M_{y}}=2.0$. (h) $\frac{M_{z}}{M_{y}}=4.0$. The Horizontal dashed lines in (e-h) correspond to the theoretical values of 0.463, 0.547, 0.628, and 0.728, respectively. The open boundary conditions are applied in the numerical calculations. The Fermi energy is $E_{F}=0.4$ in the central region and $E_{F}=0.6$ in the lead. The transport direction is along z, and the ensembles for the average are 1200.
@@ V.1 Tight binding model of $\mathrm{Cd_{3}As_{2}}$
where $a^{\prime}_{\mathbf{r}}(a^{\prime\dagger}_{\mathbf{r}})$ represents the two-component electron annihilation(creation) operators on the site $\mathbf{r}$. Based on the parameters of the real material given in [79], we set the parameters as $C_{0}=-0.0145~{}\rm{eV}$, $C_{1}=10.59~{}\rm{eV}\AA^{2}$, $C_{2}=11.5~{}\rm{eV}\AA^{2}$, $M_{0}=0.0205~{}\rm{eV}$, $M_{1}=-18.77~{}\rm{eV}\AA^{2}$, $M_{2}=-13.5~{}\rm{eV}\AA^{2}$, the SOC strength $A=0.889~{}\rm{eV}\AA$, and the lattice constant along the three directions is $a=b=3~{}\rm{\AA},c=5~{}\rm{\AA}$. These parameters apply to both the central region and the two leads. The Fermi energy in the lead is set equal to that in the central region to ensure state alignment. Open boundary conditions are applied in the conductance calculations. The onsite disorder is added and averaged over at least 1200 ensembles to calculate $\Delta G$. In the following, $\rm{eV}$ and $\rm{\AA}$ are taken as units of energy and length, respectively, and will be omitted for simplicity hereafter.
@@ V.2 UCF versus Fermi energy along y and z transport directions
The largest system size considered is $L_{x}=L_{y}=L_{z}=20$, where $L_{\alpha}$ counts the number of unit cells along the $\alpha$ direction. Unlike in previous sections, the lattice constants are not unity. The system is thus spatially anisotropic with lengths $\hat{L}_{x}=\hat{L}_{y}=60~{}\rm{\AA}$ and $\hat{L}_{z}=100~{}\rm{\AA}$, where $\hat{L}_{\alpha}$ measures the sample lengths in units of $\rm{\AA}$ along the $\alpha$ direction. As discussed at the end of Section  IV.2, this geometry will favor a significant Fermi energy dependence of UCF for transport along the $x$ and $y$ directions by decreasing the diffusion time $t_{x/y}$ in Eq. 9 and will be demonstrated below.
@@ VI Discussions and Conclusions
Motivated by recent experimental observations [51] and theoretical progress [15], this work comprehensively studies UCF in materials with anisotropic band dispersion. This introduces a significant Fermi-energy dependence of UCF, in contrast to the Altshuler-Lee-Stone theory [1, 2, 7], whose quantitative result suggests a Fermi-energy independence.
@@
As an application, we calculate the conductance fluctuations in $\mathrm{Cd_{3}As_{2}}$ using first-principle parameters near the Dirac point. We consider a specific sample setup cleaved with the same number of unit cells along the $L_{\alpha}$($\alpha=x,y,z$) direction. It is found that the UCF increases when the chemical potential moves away from the Dirac point for transport along the $x$ or $y$ direction. In contrast, for transport along the $z$ directions, the UCF dependence on Fermi energy is much less sensitive above the Dirac point. The distinct behaviors of UCFs are testable in experiments. The scope of this work addresses the Fermi energy dependence of UCF in anisotropic materials. Our work does not consider the effects of the magnetic field [86], the long-range impurities [38], decoherence [50, 87, 88, 57, 46], finite temperature [89, 90], the changes of symmetry indices [39, 40, 54], disorder-induced phase transitions [55, 30], non-ergodicity [56], etc, which will be left for future research. In addition to $\mathrm{Cd_{3}As_{2}}$, anisotropic materials are prevalent [73] in nature and can be engineered by straining [91, 92] or twisting  [93]. With the well-developed techniques of tuning the chemical potential through gate voltage or doping, we expect our findings to be readily testable in future experiments.
@@ Acknowledgements
Data Availability
The data supporting the findings of this article are available [94], embargo periods may apply.
@@ Appendix B The Fermi surface and density of states
Appendix C Numerical results of UCF for the anisotropic free electron gas on a 3D lattice
In Section. III.2, the numerical results of UCF have been compared with theoretical predictions in 2D lattices. This section supplements the discussion with numerical results in 3D lattices. Specifically, we consider a cubic central region with $L_{x}=L_{y}=L_{z}=L$, and systematically compare the numerical universal conductance fluctuation (UCF) amplitudes with theoretical predictions for different anisotropy strengths $\frac{M_{z}}{M_{y}}$.
The numerical results confirm that for weak anisotropy of $\frac{M_{z}}{M_{y}}\leq 1$ [see Fig.11 (a) and (b)], the UCF amplitudes converge well towards the theoretical predictions as the system size increases, consistent with the 2D case. For stronger anisotropy ($\frac{M_{z}}{M_{y}}>1.0$), both theoretical and numerical UCF values increase significantly. In Fig.11 (c) and (d), the numerical UCF amplitudes become slightly larger than the theoretical values for the largest size $L=25$ considered here. This discrepancy is attributed to the sensitivity of UCF to the Fermi energy in anisotropic systems, and can be mitigated by positioning the Fermi level closer to the band bottom. However, due to computational cost constraints, we do not extend the calculations to larger system sizes. Nevertheless, the observed trends provide clear evidence supporting the theoretical dependence of UCF on anisotropy in 3D systems.
Figure 11: (Color online) Numerical results of the conductance fluctuations $\Delta G$ versus the disorder strength W for different anisotropy parameters $\frac{M_{z}}{M_{y}}$ in the 3D cubic system with $L_{x}=L_{y}=L_{z}=L$ and with $M_{z}=0.5$, $M_{x}=M_{y}$. (a) $\frac{M_{z}}{M_{y}}=0.5$. (b) $\frac{M_{z}}{M_{y}}=1.0$. (c) $\frac{M_{z}}{M_{y}}=2.0$. (d) $\frac{M_{z}}{M_{y}}=5.0$. The Horizontal dashed lines in (a-d) correspond to the theoretical values of $0.458$, $0.548$, $0.618$, and $0.742$, respectively. For each $\frac{M_{z}}{M_{y}}$, the Fermi energy in the central region and lead are $E_{F}=\Delta$ at the band bottom and $E_{F}^{lead}=8\Delta$ at the band center with $\Delta=(M_{x}+M_{y}+M_{z})/4$. Several system sizes $L=10,15,20$ and $25$ are calculated to show the convergence in the thermodynamic limit. The open boundary conditions are applied in the numerical calculations. The transport direction is along z, and the ensembles for the average are 1200.